40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 343 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  360   Wed Mar 5 12:51:48 2008 JohnSummaryLSCInitial Ligo Arm finesse versus lambda
I've taken the coating recipes for the initial ligo arm cavity from Rana's web page (ligo.caltech/edu/~rana/mat/)
and plotted the finesse as a function of wavelength. There is some uncertainty over the indices of refraction but
the main conclusion remains unchanged - i.e. it appears that using other wavelengths will be difficult.
Stefan is looking at how to tune the layers of any new mirrors to make dichroic optics.
Attachment 1: FofLambdaLIGOI.jpg
FofLambdaLIGOI.jpg
  359   Wed Mar 5 12:35:09 2008 JohnSummaryComputer Scripts / ProgramsPlot photodiode responses in MatLab
A matlab function to plot the responses of photodiodes. There's still plenty of room for improvement but it should work for all our diodes without any changes. You may want to adjust which points are used in the fit to remove time delay.




% Plot data from diode response measurements
function out = diodeplot(f_Hz,mag_dB,phase_deg,f_beat_MHz)

% $$$ clear all
% $$$ close all
% $$$ clc
% $$$
% $$$
% $$$ mag = dlmread('D:\40m\PD6\M7.txt','\t', 15, 0);
% $$$ phase = dlmread('D:\40m\PD6\P7.txt','\t', 15, 0);
% $$$
% $$$ % Frequency i.e. x-axis
% $$$ f = mag(:,1);
% $$$
% $$$ % Magnitude in dB
% $$$ mag_dB = mag(:,2);
% $$$
% $$$ % Phase in degrees
% $$$ phase_deg = phase(:,2);
% $$$
% $$$ % Frequencies of interest
% $$$ f_beat_MHz = [33 133 166 199]*1e6;
% $$$
% $$$ diodeplot(f, mag_dB, phase_deg, f_beat_MHz)

% x axis limits
xmin = 10e6;
xmax = 500e6;

% Unwrap phase
phase_deg = (180/pi)*unwrap((pi/180)*phase_deg);

%Find values at our freqeuncies of interest
Mag_f_beat = interp1(f_Hz,mag_dB,f_beat_MHz);

% Remove the time delay from the phase data
% (May want to adjust which points are selected here)

straight = @(a, x) a(1) * x + a(2);

xdata = f_Hz;
ydata = phase_deg;

aguess = [10 0.1];
a = lsqcurvefit(straight,aguess,xdata,ydata);
fit = straight(a,xdata);

phase_deg = phase_deg-fit;

figure(1)
ha = axes('units','normalized','position',[0 0 1 1]);
uistack(ha,'bottom');
I=imread('PDbw.jpg');
hi = imagesc(I);
colormap gray
set(ha,'handlevisibility','off', ...
'visible','off')
plot(xdata,ydata,'r')
hold on
plot(xdata,fit,'k')
plot(xdata,phase_deg,'b')
hold off
ylabel('Phase/ degrees', 'FontSize',12)
xlabel('Frequency/ Hz', 'FontSize',12)
title('Removing the time delay','FontSize',16)
legend('data','fit','data-fit',0)
set(hi,'alphadata',.35)
set(gca,'Color','None')
box off




figure(2)
set(gcf,'Color', [1 1 1])
subplot(4,1,[1 3])

semilogx(f_Hz,mag_dB,'k','LineWidth',2.5)
title('Response of PD6','FontSize',16)
ylabel('Magnitude/ dB', 'FontSize',12)
xlim([xmin xmax])
grid

MagLayout = get(gca, 'Position');
YLimits = get(gca, 'YLim') ;
LabelExt = [];

for ivar = 1:length(f_beat_MHz);

a = text(f_beat_MHz(ivar),1.05 * min(Mag_f_beat),...
sprintf('%2.1fdB @ %dMHz', Mag_f_beat(ivar),f_beat_MHz(ivar)/1e6),...
'FontSize',10,'FontWeight','Bold','HorizontalAlignment','right',...
'VerticalAlignment','top','BackgroundColor',[.7 .9 .7],...
'Margin',0.5, 'Rotation',90);
LabelExt = [LabelExt; get(a,'Extent')];
LabelPos = get(a,'Position');

end

% Change YLim so that it is around the bottom of the labels
% There must be a better way
set(gca, 'YLim', [min(LabelExt(:,2)) YLimits(2)])

% Remove the last tick mark so that it doesn't overlap with the
% +180 of the phase plot
YTickLabelNew = str2num(get(gca, 'YTickLabel'));
YTickNew =[[] YTickLabelNew(2:end) ];
set(gca,'XTickLabel', [], 'YTick', YTickNew)

% Add lines now we know what the YLims are
for ivar = 1:length(f_beat_MHz);
line([f_beat_MHz(ivar) f_beat_MHz(ivar)], get(gca, 'YLim'))
end

subplot(4,1,4)
semilogx(f_Hz,phase_deg,'r','LineWidth',2.5)
xlim([xmin xmax])
ylabel('Phase/ degrees', 'FontSize',12)
xlabel('Frequency/ Hz', 'FontSize',16)
grid
PhaseLayout = get(gca, 'Position');
PhaseLayout(4) = MagLayout(2)-PhaseLayout(2);

% Make the top of the phase plot align to the bottom of the
% magnitude plot
set(gca, 'Color', 'None', 'Position',PhaseLayout, 'YTick',[-180:45: ...
180])
set(gcf,'units','normalized','outerposition',[0 0 1 1]);
Attachment 3: PDbw.jpg
PDbw.jpg
  358   Tue Mar 4 23:22:32 2008 robDAQComputersc1susvme1&2 rebooted

I found that some channels from c1susvme1 and c1susvme2 were not being recording by the DAQ (and were not showing up in DV). I rebooted these processors, which fix the problem. If you see other cases of this (signal exactly zero, but not a testpoint problem), just reboot the corresponding processor.
  357   Tue Mar 4 20:14:02 2008 ranaConfigurationIOOMC Alignment
The MC alignment was pretty far off. We were getting TEM01 mode locks only.
Rather than inspect what happened I just aligned the MC suspensions to get
the transmission higher. Now Matt should be able to lock the X arm and collect
adaptive filter data.
  356   Tue Mar 4 19:14:09 2008 ranaConfigurationCDSTDS & SVN
Matt, Rob, Rana

Today we added the TDS software to the 40m SVN repo.

First we rationalized things by deleting all the old TDS directories and taking
the tds_mevans dir and making it be the main one (apps/linux/tds).

We also deleted all of the TDS directories in the project area. It is now very
likely that several scripts will not work.
We're going to have the teething
problems of repointing everything to the nominal paths (in the apps areas).

Finally we did:
svn import tds https://40m.ligo.caltech.edu/svn/40m/tds --username rana

to stick it in. To check it out do:
svn checkout https://40m.ligo.caltech.edu/svn/40m/tds --username rana

We'll get a couple of the O'Reilly SVN books as well to supplement our verion control knowledge.
Unitl then you can use the SVN cheat sheets available at:
http://www.digilife.be/quickreferences/quickrefs.htm
  355   Tue Mar 4 10:08:21 2008 robUpdateComputersgreen lights unreliable when c0daqctrl down

So far I've tried powering off the framebuilder, power-cycling the RAID (it was showing an error message about bad IDE channel #4), and rebooting the LSC (just for fun). When I reset the LSC, its green light on the RFM_NETWORK screen did not turn red, making all these lights suspect. The iscepics40m process is what controls these red/green lights, so maybe it's gone wonky. It appears to be running however, on c1dcuepics, and it also seems to be functioning correctly in other ways (it's communicating correctly with the LSC).

Update: Alex and Jay came by. The solution was to reset the c0daqctrl processor, which apparently was not done in Rana's rebooting spree. Or maybe it needed to be done last.
  354   Tue Mar 4 00:42:51 2008 ranaUpdateComputersFB0 still down ?
The framebuilder is still down. I tried restarting the daqd task and resetting the RFM
switch like it says in the Wiki but it still doesn't work right. The computer itself is
running (I can ssh to it) and the daqd process is running but there's a red light for
it on the RFM screen and dataviewer won't connect to it.

If Alex isn't over by ~10 AM, we should call him and ask for help.
  353   Mon Mar 3 19:34:40 2008 ranaUpdateComputers RFM Network are down

Quote:
The CODAQ_RFNETWORK are down, except C1SUSVME & AWG

All of the FE machines were found to be down this afternoon. I called Alex and he suggested several
things which didn't work (restart EPICS tasks, power cycle RFM switch, etc.).

Then he suggested that I go around and power cycle every crate!!! And that sometimes the order of this
matters!!! I think he was just recording this conversation so that he could have a laugh by playing it on Youtube.

However, power cycling all of the FE crates seemed to work. Alex's theory is that 'something goes bad in the
RFM cards sometimes'.

Its all green now.
  352   Mon Mar 3 13:58:10 2008 steveUpdateComputers RFM Network are down
The CODAQ_RFNETWORK are down, except C1SUSVME & AWG
  351   Mon Mar 3 09:25:33 2008 steveUpdateVACrga scan logging is working now

Quote:

Quote:
The rga head and controller are running fine, but the data logging is not.


It should run tonight at 1:25 AM. To get the cron job to work properly on op340m, I had to make wrapper sh script which defines the perl library before calling the actual RGAlogger script.




Thanks to Rob, it is working !

The baked, calibrated rga head
model# M206M, s/n #c128035 was reinstalled at the 40m ifo on Feb. 8, 2008
Faraday mode dwell time was increased from 2 to 16 sec
Rga head temp at top silver gaskit 45.2 C
The noise floor is at 1E-12 Torr

There is more detail in logbook VMI-14 p 107

pd65-m-d119
Attachment 1: rgalogworks.jpg
rgalogworks.jpg
Attachment 2: cc1.jpg
cc1.jpg
  349   Sun Mar 2 23:43:45 2008 ranaHowToLSCPD6 response
John's PD plotting script is superior to all of the ones we had before; lets make him post the script so we can all use it.

Looks like PD6 is not too happy after all; the 199 MHz response is not much higher than the 166 MHz response. I thought we were supposed to have them balanced to within 6 dB or so?
  348   Fri Feb 29 13:51:17 2008 JohnSummaryLSCPD6 response
I checked the response of PD6 using the AM laser. It looks happy enough.

16 averages
-10dBm source power
77.3mV dc on the diode
  347   Thu Feb 28 19:49:21 2008 robUpdateElectronicsRF Monitor (StocMon)


Quote:

With Ben, we hooked up the RF Monitor box into the PSL rack and created 4 EPICS channels for the outputs:

C1:IOO_RF_STOC_MON_33
C1:IOO_RF_STOC_MON_133
C1:IOO_RF_STOC_MON_166
C1:IOO_RF_STOC_MON_199

The power cable bringing +15V to the preamplifier on the PSL table should be replaced eventually.


I changed the names of these channels to the more appropriate (and informative, as they're coming from the RFAMPD):

C1:IOO-RFAMPD_33MHZ
C1:IOO-RFAMPD_133MHZ
C1:IOO-RFAMPD_166MHZ
C1:IOO-RFAMPD_199MHZ

I also added them in an aesthetically sound manner to the C1IOO_LockMC.adl screen and put them in trends. Along the way, I also lost whatever Alberto had done to make these monitors read zero when there's no light on the diode. It doesn't appear to be written down anywhere, and would have been lost with a reboot anyway. We'll need a more permanent & automatable solution for this.
  346   Thu Feb 28 19:37:41 2008 robConfigurationComputersmultiple cameras running and seisBLRMS

Quote:
1) Mafalda is now connected via an orange Cat5E ethernet cord to the gigabit ethernet switch in rack in the office space. It has been labeled at both ends with "mafalda".

2) Both the GC650M camera (from MIT) and the GC750M are working. I can run the sampleviewer code and get images simultaneously. Unforutnately, the fps on both cameras seems to drop roughly in half (not an exact measurement) when displaying both simultaneously at full resolution.

3)Discovered the Gigabit ethernet card in Mafalda doesn't support jumbo packets (packets of up to 9k bytes), which is what they recommend for optimum speed.

4)However, connecting the cameras through only gigabit switches to Mafalda did seem to increase the data rate anyways, roughly by a factor of 2. (Used to take about 80 seconds to get 1000 frames saved, now takes roughly 40 seconds).

5)Need to determine the bottleneck on the cameras. It may be the ethernet card, although its possible to connect multiple gigabit cards to single computer (depending on the number of PCI slots it has). Given the ethernet cards are cheap ($300 for 20) compared to even a single camera (~$800-1500), it might be worth while outfitting a computer with multiple.


I found the SampleViewer running and displaying images from the two cameras. This kept mafalda's network so busy that the seisBLRMS program fell behind by a half-hour from its nominal delay (so 45 minutes instead of 12), and was probably getting steadily further behind. I killed the SampleViewer display on linux2, and seisBLRMS is catching up.
  345   Thu Feb 28 16:19:37 2008 josephbConfigurationComputersMafalda rewired and multiple cameras running
1) Mafalda is now connected via an orange Cat5E ethernet cord to the gigabit ethernet switch in rack in the office space. It has been labeled at both ends with "mafalda".

2) Both the GC650M camera (from MIT) and the GC750M are working. I can run the sampleviewer code and get images simultaneously. Unforutnately, the fps on both cameras seems to drop roughly in half (not an exact measurement) when displaying both simultaneously at full resolution.

3)Discovered the Gigabit ethernet card in Mafalda doesn't support jumbo packets (packets of up to 9k bytes), which is what they recommend for optimum speed.

4)However, connecting the cameras through only gigabit switches to Mafalda did seem to increase the data rate anyways, roughly by a factor of 2. (Used to take about 80 seconds to get 1000 frames saved, now takes roughly 40 seconds).

5)Need to determine the bottleneck on the cameras. It may be the ethernet card, although its possible to connect multiple gigabit cards to single computer (depending on the number of PCI slots it has). Given the ethernet cards are cheap ($300 for 20) compared to even a single camera (~$800-1500), it might be worth while outfitting a computer with multiple.
  344   Thu Feb 28 13:04:59 2008 robUpdateVACrga logging working again

Quote:
The rga head and controller are running fine, but the data logging is not.


It should run tonight at 1:25 AM. To get the cron job to work properly on op340m, I had to make wrapper sh script which defines the perl library before calling the actual RGAlogger script.
  343   Thu Feb 28 12:31:33 2008 robBureaucracyComputer Scripts / ProgramsMDV library does not work at "LINUX 2"

Quote:

While working on Thursday evening with Matlab scripts "dttfft2" and "get_data", I noticed that mDV library does not work at computer "LINUX 2" (the third computer in the control-room if you enter it from the restroom). There are multiple error messages if we try to run "hello_world", "dttfft2" or "get_data". In order to take data from accelerometers, I changed the computer - I was working from "LINUX 3" computer, the most right computer in the control room, but for the future someone should resolve the issue at "LINUX 2". I am not experienced enough to revive the correct work of mDV directory at "LINUX 2".

Andrey.


This turned out to be due to /frames not being mounted on linux2, as a result of a reboot. The issue is discussed in entry 270. I remounted the /frames, and added a line to mdv_config.m to check whether the frames are mounted.
  342   Wed Feb 27 22:05:03 2008 JohnUpdateLSCAuxiliary locking
A summary of the status of the auxiliary arm locking effort.

To help with lock acquisition we are attempting to independently lock the Y arm using light injected through ETMY. At present this secondary light source is an NPRO laser situated on the SP table. The laser light is transported to the ETM using a single mode optical fibre. In the future we might pick off some PSL light and apply a frequency shift.

We have been able to successfully mode match the fibre beam into the cavity and have been attempting lock the cavity using standard PDH signals (phase modulation sidebands are added to the light before it enters the fibre).

As yet no acceptable error signals have been produced. The demodulated RF signal is showing a time varying, bipolar dc offset.

We have minimised the residual amplitude modulation of the EOM but we expect small signals due to the undercoupled nature of the system, it could be that whatever RFAM still present is varying with time and causing this behaviour. We are also able to produce similar offsets by stressing (i.e. bending, shaking) the fibre. Could it be that the fibre is somehow converting PM into AM? Are we seeing etalon effects in the fibre or elsewhere?

If we cannot make any further progress with the existing setup we shall move the NPRO to the ETM table and try again. We are also looking into purchasing some other types of fibre.

Other things to consider are injecting through POY or using some other wavelength - neither seems obviously better.

Fiber, behavior
  341   Tue Feb 26 20:24:04 2008 AndreySummaryTMISorrow
As for that plot of three-dimensional surface, I indeed was wrong with the axis "Q_ETMX-Q_ITMX" (I put there wrong string "Q_ITMX-Q_ETMX"). On Friday plot there were values 10^(-12) on the z-axis, and that should be really meters, but the point that as I realized on Monday, I have never calibrated experimental measurement results from counts to meters , that's why it is this difference between 10^(-6) and 10^(-12). I still did not find the way to compare experim. and theoretical plots, because even if I leave "counts" on both plots, so that I have scale 10^(-6) on both plots, then the change in theoretical plot is just 0.02*10^(-6) for the range of Q-factors change, while the change in experimental measurements is an order of magnitude more 0.4*10^(-6), so the surface for theretical plot would be almost flat in the same axes as experimental results.
  340   Sun Feb 24 10:51:58 2008 tfFrogsEnvironment40m in phdcomics?
  339   Fri Feb 22 21:19:38 2008 AndreyBureaucracyComputer Scripts / ProgramsMDV library does not work at "LINUX 2"

While working on Thursday evening with Matlab scripts "dttfft2" and "get_data", I noticed that mDV library does not work at computer "LINUX 2" (the third computer in the control-room if you enter it from the restroom). There are multiple error messages if we try to run "hello_world", "dttfft2" or "get_data". In order to take data from accelerometers, I changed the computer - I was working from "LINUX 3" computer, the most right computer in the control room, but for the future someone should resolve the issue at "LINUX 2". I am not experienced enough to revive the correct work of mDV directory at "LINUX 2".

Andrey.
  338   Fri Feb 22 20:42:44 2008 AndreySummaryComputer Scripts / ProgramsIt seems I succeeded in theoretical simulations

I am pretty happy at this moment.

I definitely feel that it took me too much time to understand how to do the Matlab program and how to overcome difficulties,

but eventually at last my Matlab program seems to start working.

Briefly: What the program does?
--> take time-domain signal from two accelerometers near ITMX and ETMX (use 'get_data');
--> calculate the time-evolution of those two signals through the system "stack + pendulum" to the test-masses ITMX and ETMS (use 'lsim' in Matlab),
which gives us the time-domain evolution of the deviation of the position of individual test-mass from its average position.
--> Subtract the two results from each other in time-domain, this gives us the deviation of the length of the XARM-cavity from its average value
(roughly speaking, deviation of the length of the cavity from exactly 40 meters, although I am aware that the exact average length of XARM is less than 40 meters).
--> Take the amplitude spectrum of the result, using Sqrt(pwelch) and calibrate it from "counts" to "meters".
--> Calculate root-mean-square of peaks at different frequency intervals, for example near 0.8Hz,
and plot the three-dimensional surface showing the dependence of RMS on Q-factors Q_{ETMX} and Q_{ITMX}.

Eventually I am able to create these dependences of RMS.

I see that the minimum of the dependence is close to the diagonal corresponding to exact equality of Q_ETMX} and Q_{ITMX}, but not exactly along the diagonal. The plot allows to say
which of two conditions "Q_I > Q_E" or "Q_E < Q_I" should be fullfilled for optimization reasons. My plot is raw, I might have made a mistake in axis-label, I do not garantee now that the axis label "Q_ITMX - Q_ETMX" is correct,
maybe I need to change it for "Q_ETMX - Q_ITMX". I need some more time to determine this on Monday, but clearly there is asymmetry between Q_I and Q_E.

The peak at 0.8 Hz is pretty stable, while the peak at around 3Hz is not very repeatable, therefore in both experimental measurements and these simulations the amplitude of RMS of peak at 3Hz) is several orders of magnitude smaller than for RMS of peak at 0.8Hz, and I do not see minimum somewhere in the RMS-dependence, I see now only steady growth of RMS as Q_factors increase.

I will need to spend some time on Monday trying to understand how the sampling frequency and number of fft-points influence my results when I take amplitude spectrum using pwelch-command, as well I will need to double-check the correctness of normalization from counts to meters (I am not confident right now that amplitude of order of 10^(-12) meters is correct).

So, I need some time after the weekend to analyze my results and maybe make some slight changes, but I am glad that my Matlab model started to work in principle. I wanted to let others know about the status of the progress in my work. The fact that Matlab program works now is a good ending of a week.

Andrey.
Attachment 1: RMS_peak_08Hz-Theoretical.png
RMS_peak_08Hz-Theoretical.png
Attachment 2: RMS_peak_08Hz-QI-QE.png
RMS_peak_08Hz-QI-QE.png
Attachment 3: RMS_peak_3Hz-Theoretical.png
RMS_peak_3Hz-Theoretical.png
Attachment 4: RMS_peak_broad-interv-Theoretical.png
RMS_peak_broad-interv-Theoretical.png
  337   Fri Feb 22 16:47:54 2008 robUpdateElectronicsBaloney
Well I guess Rana didn't study too hard at Professor School, either. If he'd even bothered to actually read John's entry, he might have looked at the RF Out from the HP Analyzer. As it is, this experience so far has been like taking your car to a highly respected mechanic, telling him it's having acceleration problems, and then he takes a rag and wipes some dirt off the hood and then tells you "It's running fine. That'll be 500 bucks."

I make the current score:

Snarkiness: 2
Education:  0



I did RTFM, and it doesn't mention anything about crazy behaviour on the RF Output. So, I set up the analyzer to do a sweep from 500MHz to 1MHz, with output power of 0dBm, and plugged the output directly into the 300MHz scope with the input set to 50 Ohm impedance. The swept sine output looks totally normal from 500Mhz to 150MHz (measuring ~220mVrms below 300MHz -- 0dBm), where it abruptly transitions to a distorted waveform which the scope measures as having a frequency of ~25MHz and with 450mVrms (+6dBm). It then transitions again at some other part of the sweep to a cleaner-looking 25MHz waveform with ~1.2Vrms (+15dBm). See the attached Quicktime movie. The screeching in the background is the PSL door.

With this bizarre behaviour, it's actually possible that even someone who does everything carefully and correctly could break sensitive electronics with this machine. Let's get it fixed or get a new one.

Don't use the HP4195A anymore unless you want broken electronics.


Quote:
I'm not sure where Ward and Miller went to Analyzer school, but it was probably uncredited.
I turned it on and used 2 BNC cables and a T to hook up the source to the 2 inputs and measured the always-exciting TF of cable.

Score:  HP Analyzer  1
        Rob & John   0


I have left the analyzer on in this complicated configuration. RTFM boys.


Quote:
The HP 4195A network analyser may be broken, measurements below 150MHz are not reliable. Above 150MHz everything looks normal. This may be caused by a problem with its output (the one you'd use as an excitation) which is varying in amplitude in a strange way.

Analyzer
Attachment 1: bustedHP.mov
  336   Fri Feb 22 15:16:33 2008 AndreyUpdatePEMITMX Accelerometer is NOT broken

As I wrote in message 330, there was a bad signal from ITMX accelerometer. I have found the reason: the BNC-cable which goes from the black board with switches for accelerometer gain (1,10,100) towards DAQ-tower was completely disconnected from that black board with gain-switches. The end of the long BNC-cable was on the floor. Therefore, it was totally impossible to see any accelerometer signal. The cable that I am writing about should transport the signal from ITMX_X accelerometer.

Now all the BNC-connections seem to be in good shape, and spectra of accelerometers near ITMX and ETMX , both of them are in x-directions, are very much similar.
Attachment 1: Accelerom-ITMX-Feb23.jpg
Accelerom-ITMX-Feb23.jpg
Attachment 2: Accelerom-ETMX-Feb23.jpg
Accelerom-ETMX-Feb23.jpg
  335   Fri Feb 22 14:45:06 2008 steveUpdateMOPAlaser power levels

At the beginning of this 1000 days plot shows the laser that was running at 22C head temp
and it was send to LLO

The laser from LHO PA#102 with NPRO#206 were installed at Nov. 29, 2005 @ 49,943 hrs
Now,almost 20,000 hrs later we have 50% less PSL-126MOPA_AMPMON power
Attachment 1: lpower1000d.jpg
lpower1000d.jpg
  334   Fri Feb 22 11:13:15 2008 robUpdateElectronicsRF Monitor (StocMon)

Quote:
It took some times because the written procedure to start the chiller is not very precise.


It is actually very precise. Precisely wrong.
  333   Fri Feb 22 11:11:00 2008 robUpdateElectronicsREFLDD problem found

I used a network analyzer that actually works to find a problem in the REFLDD electronics chain. There was loose (=bad) SMA-BNC adaptor on the output of channel one of the HP RF Amplifier. It worked intermittently, so going onto the ISCT and fiddling with cables could sometimes temporarily fix the problem. The bad adaptor has been given to Andrey to discard.
  332   Fri Feb 22 08:33:18 2008 AlbertoUpdateElectronicsRF Monitor (StocMon)

Quote:
I put the amplifiers next to the monitor on the PSL table, layed the power and the RF SMA cables out to the rack. I'm powering the box and the amplifiers with the power supply, waiting for someone to show me tomorrow how to connect it to the Sorensen (Steve, Ben?).

I'm ready to hook up the channels into EPICS.


With Ben, we hooked up the RF Monitor box into the PSL rack and created 4 EPICS channels for the outputs:

C1:IOO_RF_STOC_MON_33
C1:IOO_RF_STOC_MON_133
C1:IOO_RF_STOC_MON_166
C1:IOO_RF_STOC_MON_199

The power cable bringing +15V to the preamplifier on the PSL table should be replaced eventually.
  331   Fri Feb 22 08:29:07 2008 AlbertoUpdateElectronicsRF Monitor (StocMon)

Quote:
I put the amplifiers next to the monitor on the PSL table, layed the power and the RF SMA cables out to the rack. I'm powering the box and the amplifiers with the power supply, waiting for someone to show me tomorrow how to connect it to the Sorensen (Steve, Ben?).

I'm ready to hook up the channels into EPICS.


Me and Ben Abbot were plugging the cables that power that RF Monitor box into the PSL rack when inadvertently we made some arcs spark between the pins on the back of one of the ADC. Somehow that made the laser shut down although the MOPA stayed on. We also notice some smell of burn.

Later on, after several failed attempts, Rob, Ben and Steve could restart the laser. It took some times because the written procedure to start the chiller is not very precise.
  330   Fri Feb 22 02:51:20 2008 AndreyUpdatePEMAccelerometer ITMX seems to be broken

As people probably know,

I am trying (for a long time) to create a computational program that calculates the evolution of accelerometer time-domain data through stacks and pendulum transfer functions to test masses, and calculate the RMS of differential arm lenght spectrum.

I noticed on Tuesday that time-domain signals from the two accelerometers (one is near ETMX, the other one is near ITMX) seem to have different amplitudes of fluctuations around the mean value. I suspected that this is the main reason why I cannot get the awaited result of minimum of RMS for equal values of Q-factors for ETMX and ITMX suspensions (because we subtract two very different numbers, so we cannot get anything close to zero). I took amplitude spectra of the accelerometer data (dttfft2), and they look very differently for ETMX and ITMX accelerometers. I believe that spectrum of ETMX accelerometer represents seismic noise, but accelerometer ITMX seems to provide us with irrelevant and wrong data. No peaks, just almost monotoneous decreasing curve, and 10 times smaller amplitude. Therefore, ITMX seems to be broken.

I will try tomorrow to clap my hands, shout, yell, near the broken accelerometer to confirm that the accelerometer is broken (more precisely, that either accelerometer itself is broken,
or cable connections, or DAQ channel, but something is wrong). Now it is very late, and I am going home.

See attached figures: time-scale is 10^(-1), 10^0, 10^1, 10^2 Hz.
Attachment 1: Accelerom-EYMX-Feb22.jpg
Accelerom-EYMX-Feb22.jpg
Attachment 2: Accelerom-ITMX-Feb22.jpg
Accelerom-ITMX-Feb22.jpg
  329   Thu Feb 21 19:55:46 2008 ranaUpdateElectronics2 BNC Cables, 1 Tee
I'm not sure where Ward and Miller went to Analyzer school, but it was probably uncredited.
I turned it on and used 2 BNC cables and a T to hook up the source to the 2 inputs and measured the always-exciting TF of cable.

Score:  HP Analyzer  1
        Rob & John   0


I have left the analyzer on in this complicated configuration. RTFM boys.


Quote:
The HP 4195A network analyser may be broken, measurements below 150MHz are not reliable. Above 150MHz everything looks normal. This may be caused by a problem with its output (the one you'd use as an excitation) which is varying in amplitude in a strange way.

Analyzer
  328   Thu Feb 21 18:29:28 2008 JohnSummaryGeneralHP Network Analyser Analyzer
The HP 4195A network analyser may be broken, measurements below 150MHz are not reliable. Above 150MHz everything looks normal. This may be caused by a problem with its output (the one you'd use as an excitation) which is varying in amplitude in a strange way.

Analyzer
  327   Thu Feb 21 09:56:26 2008 steveUpdatePSLthe laser is back
The laser and the psl recovered.
The water chiller temp is 19.98C and head temp 18.4C
Power 3.2W
PMC_T 3.1

c1iovme was restarted and the mc is locking now
Attachment 1: laserisback.jpg
laserisback.jpg
  326   Wed Feb 20 16:24:37 2008 steveUpdatePSLthe laser is recovering slowly
Head temp is still 20.5C and decreasing slowly.
Power output 2.9W
NPRO power 22 mW is increasing as head is cooling down
Attachment 1: laserecoverring.jpg
laserecoverring.jpg
  325   Wed Feb 20 11:34:17 2008 steveUpdatePSLlaser head temp is up
MOPA head temp is running at 20.3C now
Nomally it is at 18.5C
Attachment 1: htempup.jpg
htempup.jpg
  324   Tue Feb 19 18:28:41 2008 JohnUpdatePEMMore seismic in Baja California
Steve spotted more activity from the same quake.

Reset watchdog on ETMY.
  323   Tue Feb 19 15:21:47 2008 AndreyUpdateSUSEarthquake tripped watchdogs in ETMY, ITMY

According to the web-page http://earthquake.usgs.gov/eqcenter/recenteqsus/Quakes/ci14351140.php ,

there was a 5.0 earthquake in northern Baja California in Mexico at 02.41PM earlier today.

This earthquake made an effect on our watchdogs for ETMY and ITMY (their currents exceeded maximal values).
Watchdogs for ITMY are now restored back,
and it is taking more time for a "side degree" for ETMY to calm down,
it is still (40 minutes after the kick) swinging a lot with amplitude ~ 200mV.
  322   Tue Feb 19 10:14:13 2008 steveUpdateVACrga logging needs help
The rga head and controller are running fine, but the data logging is not.
  321   Mon Feb 18 12:04:39 2008 AlbertoUpdateElectronicsRF Monitor (StocMon)
I put the amplifiers next to the monitor on the PSL table, layed the power and the RF SMA cables out to the rack. I'm powering the box and the amplifiers with the power supply, waiting for someone to show me tomorrow how to connect it to the Sorensen (Steve, Ben?).

I'm ready to hook up the channels into EPICS.
Attachment 1: DSC_0443.JPG
DSC_0443.JPG
  320   Fri Feb 15 22:16:04 2008 AndreyUpdateComputersMATLAB is not working: "Licence checkout failed"

For some unknown to me reason,

Matlab stopped working about 20 minutes ago on all computers in the control room (both UNIX control machines and Windows).

It says: "License checkout failed. License Manager Error -15. MATLAB is unable to connect to the license server."

I do not know how to revive Matlab.

At the same time, I consider that I made a significant progress in building my theoretical/computational model in the last 2 days. I was able to compute the time-evolution of accelerometer signals through stacks and pendulums using Matlab command "lsim", and I am now able to calculate RMS of spectrum of differential arm length in different frequency intervals. It seemed to me that everything is ready in my program to make the three-dimensional theoretical/computational plot (RMS as a function of Q-factors of ETMX and ITMX), but unfortunately Matlab stopped working. It seemed to me that all that was remaining was to run a loop with all possible values of Q-factors. Let's hope that Matlab will be working after the weekend.

Andrey.
  319   Fri Feb 15 10:28:44 2008 ranaUpdateComputersNoise budget code changes

Quote:
In cvs/cds/caltech/NB/matlab/utilities/LSCmodel.m at line 146
I have hardwired in changes to struct lsc. Please see code.

The IFOin variable (which I admit is not documented) should refer to a file called
C1IFOparams.m in the ReferenceData directory. This does not yet exist but can be
created by merging L1IFOparams.m with our knowledge of the 40m IFO.
  318   Thu Feb 14 17:21:53 2008 Max JonesUpdateComputersNoise budget code changes
In cvs/cds/caltech/NB/matlab/utilities/LSCmodel.m at line 146
I have hardwired in changes to struct lsc. Please see code.
  317   Thu Feb 14 15:05:18 2008 robDAQDMFseisBLRMS 1.1
> 
> Also, for reasons of aesthetics, I changed the data delay from 6 minutes to 5 minutes.  We'll see if that's enough.

5 minutes didn't work.  
  316   Thu Feb 14 15:04:53 2008 robDAQDMFDMF delay

Sometime ago I edited seisBLRMS to keep of track of how long it was taking to write RMS data (that is, the delay between the accelerometer data and the write of the EPICS rms data). Here's a plot of that info, showing how the delay increases over time. I think this indicates a logical flaw in the timing of the seisBLRMS program, which sort of relies on everything running well consistently; this should not be difficult to fix. I'll maybe try increasing the delay to ~10 minutes, and making it relatively inflexible.
Attachment 1: DMFdelay.png
DMFdelay.png
  315   Wed Feb 13 20:37:11 2008 JohnUpdateLSCFibre locking - Fiber
Sam and I observed fringes in the light reflected from the Y arm. These fringes are due to the sidebands and not the carrier. To improve matters we plan to reduce the RF AM and increase our modulation index.
  314   Wed Feb 13 11:41:00 2008 AlbertoUpdateElectronicsSome characterization of the RF Monitor Box (StocMon)
I'm attaching a table with some measurements and the power spectrum from the pd to help evaluate the numbers.

The box output ranges from 0.5V to 2.1V. The coefficient between power and voltage is negative so higher voltage means lower power.

The red numbers are the outputs from each channel at their resonant frequencies. As one can see these are not very well centered on the dynamic range of the power detectors.

The cross coupling seems to be not a problem.

Even if the 166 filter, which handles the smallest of the frequencies and is also the most lossy (for construction reason), mounts a preamplifier, the output is still rather small. this explain also the high bias due to the noise amplification at the maximum power (13dB). A better insertion loss either remaking the filter or re-tuning that one would simplify many problems, i.e. there is not much room in the metal pomona box to fit the amplifier. I might want to consider, after everything else is ready and if I have time before leaving next week, to work on a new 166 filter.
Attachment 1: CircuitCharacterization.png
CircuitCharacterization.png
Attachment 2: alberto.spectrum3.png
alberto.spectrum3.png
  313   Tue Feb 12 16:39:52 2008 robUpdateLockingreport

Did some locking work on DRFPMI on sunday and (with John) on monday nights. So far progress has not been terribly encouraging.

Problems include the DD_handoffs not working and the CARM->MCL handoff not working so well. To get around the DD signals trouble, I decided for now to just ignore 67% of the DD signals. We should be able to run with PRC & MICH on single demod signals, and SRC on a DD signal. This seems to work well in a DRMI state, and it also works well in a DRMI+2ARMs state.

The CARM->MCL handoff actually works, but it doesn't take kindly to the AO path and it doesn't work very stably. I guess this was always the most fragile part of the whole locking procedure, and it's fragility is really coming to light now. Investigation continues.
  312   Tue Feb 12 16:34:07 2008 robDAQDMFseisBLRMS 1.1
The compiled version of seisBLRMS had been running ~2 weeks without crashing as of last night, when I killed it
so it wouldn't interfere with alignment scripts.  I added an EPICS channel C1:DMF-ENABLE, and updated the DMF
executables to check this channel while running.  So far it seems to work.  When you're running alignment acripts,
simply click the DISABLE button on the C1DMF_MASTER.adl screen, and then re-ENABLE when the scripts finish. 
It's not clear why this is necessary though.  Theories include the constant disk access is keeping the
framebuilder busy, reducing its ability to deal with ezcademod commands and DMF programs just flooding the
network with so much traffic that ezcademod-related packets run late and get ignored.

Also, for reasons of aesthetics, I changed the data delay from 6 minutes to 5 minutes.  We'll see if that's enough.
  311   Tue Feb 12 16:18:29 2008 robConfigurationComputer Scripts / Programsautoburt cron moved to op340m

I moved the autoburt cron job from op440m to op340m. For some reason, burtrb requires gcc to run (I gather it uses the C-preprocessor to parse input files), so I had to install that on op340m to get it to work properly.

There are no more cron jobs running on op440m now.
  310   Tue Feb 12 13:53:27 2008 robOmnistructureVACReturn of the RGA

The new RGA head was installed a few days ago. I just ran the RGAlogger script to see if it works, which it does. I also edited the crontab file on op340m to run the RGAlogger script every night at 1:25 AM. It should run tonight.
Attachment 1: RGA-080212.png
RGA-080212.png
ELOG V3.1.3-